a chemist needs a 30% iron alloy. How many grams of a 70% iron alloy must be mixed with 16 grams of a 20% iron alloy to obtain the required 30% alloy
Answer by stanbon(48558)

a chemist needs a 30% iron alloy. How many grams of a 70% iron alloy must be mixed with 16 grams of a 20% iron alloy to obtain the required 30% alloy
----------------------
Equation:
iron + iron = iron
0.70x + 0.20*16 = 0.30(x+16)
Multiply thru by 100 to get:
70x + 20*16 = 30x + 30*16
40x = 10*16
x = 4 grams (amount of 70% alloy needed in the mix)
=======================================================
Cheers,
Stan H.